A vending machine dispenses coffee, and a random sample of 27 filled cups have contents with a mean of 7.14 oz and a standard deviation of .17 oz. Use a .05 significance level to test the claim that the true mean fill for all cups is greater than 7 oz. You may assume a normal population (n < 30 here; assume that assumption was checked). Fill in and show calculations.

Ho                                                       S.V.

Ha                                                       Dec.

T.S.                                                      Concl.

R.R. (draw and label value)

                                                            p-value (indicate one)   > .05   or   < .05
Answer #1

solution Test statutic Here Ho 4= 7 vs Hills7 Given n=27 x = 7.14, s=0.17 a ll - 7.14-1-4: 2792 t= - 01/07 d-f=n-1=27-1=26 ;